# Large Language Models and Code Security: A Systematic Literature Review

### v1

@@ III-C Selection Criteria
IN1: The title and abstract of the paper include the search keywords
IN2: Complete papers presented at specified conferences or published in journals, or available as preprints on arXiv
@@
IN5: Papers that explore the impact of the poisoned training data of LLMs on security and their ability to detect and fix vulnerabilities
@@
EX3: Secondary studies, such as SLRs, general reviews, or surveys
@@ IV Security Vulnerabilities introduced by LLMs
Before delving into a detailed exploration of each category, we think it is worth analyzing how often these categories are discussed, throughout the 20 papers. As shown in Figure 2, the attention that has been given by the various authors to the different categories, is fairly distributed, with a few categories deemed more critical than others. For example, injection vulnerabilities are covered in 16 out of 20 studies, while error handling vulnerabilities are discussed in only 3 studies.
@@ IV-B Memory Management
This category includes vulnerabilities related to memory management, which were also identified across most of the included studies. These vulnerabilities include overflow vulnerabilities, such as buffer and integer overflow, mostly caused by the lack of input sanitization. Moreover, this category includes null pointer dereference and use-after-free memory allocation vulnerabilities. These issues arise due to improper handling of memory references and pose significant risks to system security. Although [23, 24, 26, 27, 28, 31, 32, 11, 33, 35, 39] all identified memory vulnerabilities in code produced by LLMs, Liu et al. [23] placed special importance on memory management vulnerabilities. They performed a systematic empirical assessment of the quality of code generated using ChatGPT, by solving 728 algorithmic problems in 5 languages. Their study highlighted significant issues in memory management within LLM-generated code. Specifically, they found that the majority of vulnerabilities were related to missing null pointer tests, with 91.8% of the total vulnerabilities attributed to the MissingNullTest query. These code snippets failed to check for null after memory allocation, which can lead to critical security vulnerabilities such as CWE-476: Null pointer dereference, also identified in [27, 28, 32, 35].
@@ IV-F Authentication and Authorization
Hamer et al. [36] compared the security properties of the code generated by ChatGPT, with answers provided in Stack Overflow. They concluded that ChatGPT-generated code had 20% fewer vulnerabilities than Stack Overflow snippets, although both produced significant security flaws. One of the most frequent CWEs in both ChatGPT-generated code and Stack Overflow snippets was CWE-798: Use of Hard-coded Credentials. CWE-798 is also discussed in [23, 28, 30, 40].
@@ IV-G Cryptography
Cryptographic vulnerabilities arise when encryption mechanisms are poorly implemented or completely absent. Common causes of weak encryption include the use of broken or insecure cryptographic algorithms, improper validation of certificates, the use of weak or absent hashing functions, and the reuse or hard-coding of cryptographic keys. Several studies [24, 30, 33, 36, 37, 38, 39, 10, 40] identified cryptographic weaknesses in LLM-generated code.
@@
A common vulnerability identified in multiple studies is CWE-327: Use of a Broken or Risky Cryptographic Algorithm [33, 36, 37, 39].
@@ IV-H Resource Management
In general, resource management vulnerabilities were identified by [25, 27, 32, 37, 38, 10, 40].
@@ IV-I Coding Standards
Additionally, CWE-758: Reliance on Undefined, Unspecified, or Implementation-Defined Behavior is another significant concern discussed in by Kim et al. [32], and Sandoval et al. [27]. Besides the mentioned studies, others have focused on this type of vulnerability [25, 29, 38, 40].
@@ V-A Uses of LLMs for Vulnerability Detection
A study by Zhang et al. [46] focused on LLMs capabilities for automated vulnerability detection. They tested over 10 different LLM models using various techniques like zero-shot and one-shot learning (for more details on prompting techniques, we refer the reader to Section VIII), and fine-tuning. They applied these models to various datasets of source code, exhibiting C/C++ and smart contract vulnerabilities. The study found that CodeLlama-7B achieved the highest F1-score of 82% with discriminative fine-tuning, along with a precision of approximately 89% and a recall of around 78%. In comparison, other methods were less effective. The authors have also identified issues with input length limits, and proposed a solution based on sliding windows that divide long sequences into overlapping sequences of tokens. In addition, they introduced a right-forward embedding strategy, which enhances context by making each token aware of tokens that come after. Overall, LLMs appear to be valuable tools for vulnerability detection in real-world scenarios.
According to Wen et al. [47], existing pre-trained models for vulnerability detection often miss important details and struggle with complex code. As a consequence, performing accurate vulnerability detection can be difficult. To address these challenges, Wen et al. [47] proposed a Structured Natural Language Comment tree-based Vulnerability Detection framework, based on pre-trained models, named SCALE. This framework is composed of three main modules. The first module, Comment Tree Construction, helps the model understand code semantics, by incorporating LLMs for comment generation and adding the comment node to Abstract Syntax Trees (ASTs). The second module, Structured Natural Language Comment Tree Construction, refines this process by explicitly involving code execution sequences and combining code syntax templates with the comment tree. The third module, SCT-Enhanced Representation, uses these improved comment trees to capture vulnerability patterns and better understand vulnerabilities in the code.
Results show that SCALE outperforms existing methods, including supervised, retrained model-based, and other LLM-based approaches in detecting vulnerabilities. SCALE achieved notable improvements in precision, recall, and F1 scores across multiple datasets, achieving, for instance, up to around 65% F1 score.
@@
Li et al. [52] noticed that LLMs often produce false positives and false negatives when used as static code analyzers, so they proposed an innovative approach to address this issue. Their method prompts ChatGPT to summarize functions, breaking down tasks step by step and progressively refining results to improve accuracy. In the study, GPT-3.5 successfully eliminated 8 out of 20 false positives, while GPT-4 performed even better, removing 16 out of 20 false positives. In conclusion, their approach improves bug detection accuracy by leveraging GPT-4 enhanced capabilities, although some inconsistencies persist.
Hu et al. [53] focused on smart contract vulnerability detection. They revealed that generating more answers with higher randomness, significantly increases the likelihood of producing correct responses, but also leads to a higher number of false positives. To mitigate this issue, they proposed a framework called GPTLens, which splits the conventional one-stage detection into two phases: generation and discrimination. In this framework, an LLM plays a dual role. In the generation phase, it acts as the Auditor, aiming to generate a wide range of vulnerabilities to capture the correct ones. In the discrimination phase, it acts as the Critic, evaluating these vulnerabilities to reduce false positives.
The authors validated the effectiveness of GPTLens by testing it on 13 smart contracts with known vulnerabilities and comparing its performance against 6 competing methods, to assess how well each approach identified true vulnerabilities while minimizing false positives. Empirical results show that GPTLens significantly improves vulnerability detection over traditional methods.
LLMs often struggle to perform complex reasoning over code to detect vulnerabilities, especially since this requires analyzing an entire repository, as discussed by Li et al. [54]. To address this, they proposed IRIS, a method that combines LLMs with SATs. IRIS uses LLMs to find specific taint specifications related to a vulnerability class (CWE) in third-party library APIs within a project. IRIS improves these specifications with CodeQL to find security vulnerabilities. They evaluated IRIS on dataset they created with 120 security vulnerabilities found in real-world Java projects, and validated manually. Results showed that IRIS detects 69 out of 120 vulnerabilities using GPT-4 while also reducing the number of false positives by 80%. In summary, the combination of LLMs and SATs has enhanced the analysis of vulnerabilities.
Li et al. [55] proposed LLift, a framework that combines LLMs with static analysis (using UBITect, a UBI discovery tool), to improve bug detection in large and complex codebases, with a focus on Use Before Initialization (UBI) bugs in the Linux kernel. First, the pipeline uses UBITect to perform a path-insensitive static analysis for scalability, followed by symbolic execution for precision. Then, in the second phase of the pipeline an LLM is used to address the undecided cases that UBITect cannot handle. The evaluation results showed that LLift reduces the complexity of path-sensitive analysis, improving bug detection accuracy without a high rate of false positives. Additionally, LLift was able to identify 4 undiscovered UBI bugs in the Linux kernel that UBITect could not confirm.
@@ V-B Issues in Using LLMs for Vulnerability Detection
While LLMs show considerable promise in vulnerability detection, several challenges and limitations persist. One of the key challenges for LLMs is the occurrence of false positives, where the model incorrectly flags a non-existent vulnerability.
For example, Steenhoek et al. [56] searched for systematic optimal prompts and proposed three prompting methods, which they used for evaluating the vulnerability detection abilities of 11 LLMs. They found that, although their proposed prompting methods enhanced LLMs performance, the LLMs still struggled with vulnerability detection. Specifically, these models achieved only 50% to 63% Balanced Accuracy, often performing close to random guessing in many scenarios. These models frequently produced false positives, mistakenly identifying vulnerabilities where none existed, a problem often referred to as hallucinations.
Adding to that, Happe et al. [57] also discussed hallucinations phenomenon in LLMs. While not frequent, the authors claim that the presence of hallucinations implies the need for better grounding and context-awareness, especially when using LLMs for critical tasks, such as security testing. The study explored the use of LLMs in penetration testing, proposing their application in two key areas: high-level task planning and low-level vulnerability hunting. For high-level tasks, LLMs assist by creating attack strategies for various systems, suggesting realistic attack vectors. For low-level tasks, the LLMs interact with a vulnerable virtual machine via SSH, analyzing the machine state and suggesting specific attacks that are then automatically executed. This demonstrates that LLMs can be useful in penetration testing, but the potential for hallucinations reinforces the importance of ensuring that the model outputs are reliable and contextually grounded.
Yu et al. [58] explored LLMs potential in security code review. To do that, they compared the performance of 6 LLMs, using 5 different prompt types, against traditional SATs. The results indicate that GPT-4 outperformed other LLMs in detecting security issues, while GPT-4 Turbo, Gemini Pro, and Llama 2 70B showed lower performance. GPT-4 was particularly effective when analyzing code with fewer tokens and when provided with a CWE list for reference. However, the capabilities of state-of-the-art LLMs in conducting security code review are still limited and often demonstrate randomness, especially in the case of the best performing LLMs. Additionally, GPT-4 suffers from issues with verbosity and non-compliance, meaning it sometimes fails to meet prompt requirements or omits expected details, which can affect the quality of its responses.
Similarly, Ullah et al. [59] remarked that LLMs encounter difficulties when detecting vulnerabilities in code. They concluded this by developing an automated evaluation framework that evaluates whether LLMs can identify and reason about security-related bugs. The evaluation shows that LLMs give unpredictable responses, have incorrect and inaccurate reasoning, and perform poorly in real-world situations. Remarkably, even advanced models like PaLM2 and GPT-4 exhibit significant weaknesses, yielding incorrect answers in 26% and 17% of cases, respectively.
Purba et al. [60] also reported weak performance of LLMs in detecting software vulnerabilities, in particular SQL injection and buffer overflow vulnerabilities. They performed an experiment applying 4 well-known LLMs (GPT-3.5-Turbo, Davinci, CodeGen, and Codegen2B-multi), to two popular public datasets, with the goal of evaluating their performance in vulnerability detection tasks. The results showed that, while LLMs performed well in identifying patterns related to vulnerabilities, especially when fine-tuned, they also exhibited high false positive rates. For example, the fine-tuned Davinci model achieved a 94% recall for buffer overflow vulnerabilities, but its false positive rate was 62.67%. They reported a gap between LLMs and popular SATs, but also highlighted the potential of LLMs for identifying subtle patterns associated with software vulnerabilities.
In their recent study, Liu et al. [61] proposed VulDetectBench, a new benchmark for evaluating vulnerability detection. This benchmark is composed of 5 tasks of increasing difficulty, and evaluates LLMs ability to identify, classify, and locate vulnerabilities. Their evaluation of 17 models found that LLMs perform well (over 80% accuracy) in simple tasks, such as identifying and classifying vulnerabilities. However, LLMs achieve less than 30% accuracy when faced with more complex tasks, such as detecting the root causes of vulnerabilities across extensive codebases. In conclusion, even though LLMs can be useful for vulnerability analysis, they still cannot replace traditional tools for more complex tasks.
Ding et al. [62] identified flaws in existing vulnerability datasets, including poor data quality, low label accuracy, and high levels of duplication. These flaws can lead to unreliable model performance, when datasets are utilized in realistic vulnerability detection scenarios. Besides, the authors argue that the evaluation methods used with these datasets are not representative of real-world vulnerability detection. To address these issues, they developed a new vulnerability dataset named PrimeVul, and introduced new evaluation guidelines to enhance the rigor of model assessment. They evaluated LLMs on PrimeVul and found that other benchmarks largely overestimate the performance of these models. For example, a state-of-the-art 7B model scored 68.26% F1 score on BigVul, but only 3.09% F1 score on PrimeVul. Overall, they concluded that LLMs are not able to meet the requirements of effective vulnerability detection in real-world scenarios.
Çetin et al. [63] evaluated LLMs ability to identify vulnerabilities in PHP code, addressing also the issue of high false positive rates. They observed that GPT-4 had the highest reliability in detecting vulnerabilities, ChatGPT-3.5 and Claude showed lower performance, and some models (e.g., Bard and Llama-2) had notably lower detection rates. Regardless of their performance, all LLMs exhibited high false positive rates, with GPT-4, the top-performing model, having close to 63% of false positives. Other LLMs performed even worse, with false positive rates reaching up to 97%. The authors also examined the performance obtained by combining different LLMs. Although the results were better than individual models, the improvements were not remarkable. Despite these results, the authors concluded that LLMs, with additional fine-tuning, can still be effective tools for static analysis.
Pelofske et al. [64] reached a similar conclusion, finding that LLMs are not reliable enough for fully automated vulnerability detection. Their study focused on evaluating the effectiveness of 5 GPT models for vulnerability detection in C/C++ code. In the best case, the recall and precision reached 100%, showing high accuracy for certain vulnerability types. This was achieved using Llama-2-70b-chat-hf GPT model at a temperature of 0.1. However, despite this success, some models still exhibited significant rates of false positives and false negatives, making them unsuitable for practical use.
Chen et al. [65] focused on ChatGPT performance in identifying vulnerabilities in blockchain smart contracts. While ChatGPT was able to identify some vulnerabilities, it struggled with precision, resulting in false positives. Additionally, approximately 20% of ChatGPT conclusions were uncertain, and the maximum code sizes that GPT-4 and GPT-3.5 models could effectively analyze were around 33 KB and 16 KB, respectively. In conclusion, while ChatGPT has potential as a vulnerability detection tool, improvements are necessary for its broader adoption in blockchain security.
Similar to Chen et al. [65], Sun et al. [66] noted shortcomings in GPT performance on smart contract vulnerabilities. To remedy, they proposed a tool named GPTScan, which combines GPT with static analysis techniques to address deficiencies in detecting logic vulnerabilities. GPTScan first parses the contract code, and uses GPT to identify potential vulnerability scenarios and key variables. Then, GPTScan matches these potential vulnerabilities to known vulnerability types, and passes the results to a static analysis module for confirmation. The results indicate that GPTScan can achieve a precision of over 90% for token contracts, a 57.14% precision for large projects, and a recall of over 70% for detection of ground-truth logic vulnerabilities. GPTScan is fast, cost-effective, and capable of discovering new vulnerabilities missed by humans.
Last, Gao et al. [67] conducted a large-scale study to measure the performance of 16 LLMs in vulnerability detection by means of VulBench, a benchmark proposed by the authors to address existing dataset quality issues. They highlighted the potential of LLMs in this area, particularly GPT-4, which outperformed other models in vulnerability detection tasks, on relatively simple datasets. However, the results have shown that performance drops across models, when LLMs face complex, real-world, detection tasks.
V-C Comparison with Other Approaches

### v2

@@ III-C Selection Criteria
IN1: The title and abstract of the paper include the keywords.
IN2: Complete papers presented at specified conferences or published in journals, or available as preprints on arXiv.
@@
IN5: Papers that explore the impact of the poisoned training data of LLMs on security and their ability to detect and fix vulnerabilities.
@@
EX3: Secondary studies, such as SLRs or surveys.
@@ IV Security Vulnerabilities introduced by LLMs
Before delving into a detailed exploration of each category, we think it is worth analyzing how often these categories are discussed, throughout the 20 papers. As shown in Figure 2, the attention that has been given by the various authors to the different categories, is fairly distributed, with a few categories deemed more critical than others. For example, injection vulnerabilities are covered in 16 out of 20 studies, while error handling vulnerabilities are discussed in 3 studies.
@@ IV-B Memory Management
This category includes vulnerabilities related to memory management, identified across most of the included studies. These vulnerabilities include overflow vulnerabilities, such as buffer and integer overflow, mostly caused by the lack of input sanitization. Moreover, this category includes null pointer dereference and use-after-free memory allocation vulnerabilities. These issues arise due to improper handling of memory references and pose significant risks to system security. Although [23, 24, 26, 27, 28, 31, 32, 11, 33, 35, 39] all identified memory vulnerabilities in code produced by LLMs, Liu et al. [23] placed special importance on memory management vulnerabilities. They performed a systematic empirical assessment of the quality of code generated using ChatGPT, by solving 728 algorithmic problems in 5 languages. Their study highlighted significant issues in memory management within LLM-generated code. Specifically, they found that the majority of vulnerabilities were related to missing null pointer tests, with 91.8% of the total vulnerabilities attributed to the MissingNullTest query. These code snippets failed to check for null after memory allocation, which can lead to critical security vulnerabilities such as CWE-476: Null pointer dereference, also identified in [27, 28, 32, 35].
@@ IV-F Authentication and Authorization
Hamer et al. [36] compared the security properties of the code generated by ChatGPT, with answers provided in Stack Overflow. They concluded that ChatGPT-generated code had 20% fewer vulnerabilities than Stack Overflow snippets, although both produced security flaws. One of the most frequent CWES in both ChatGPT-generated code and Stack Overflow snippets was CWE-798: Use of Hard-coded Credentials. CWE-798 is also discussed in [23, 28, 30, 40].
@@ IV-G Cryptography
Cryptographic vulnerabilities arise when encryption mechanisms are poorly implemented or completely absent. Common causes of weak encryption include the use of broken or insecure cryptographic algorithms, improper validation of certificates, the use of weak or absent hashing functions, and the reuse or hard-coding of cryptographic keys. Several studies [24, 30, 33, 36, 37, 38, 39, 10, 40] identified cryptographic weaknesses in LLM-generated code, without exploring them in detail.
@@
A common cryptographic vulnerability identified in multiple studies is CWE-327: Use of a Broken or Risky Cryptographic Algorithm [33, 36, 37, 39].
@@ IV-H Resource Management
In general, resource management vulnerabilities were identified by [25, 27, 32, 37, 38, 10, 40], however, none of these studies explored in depth the resource management vulnerabilities in LLM generated-code.
@@ IV-I Coding Standards
Additionally, CWE-758: Reliance on Undefined, Unspecified, or Implementation-Defined Behavior is another significant concern discussed in by Kim et al. [32], and Sandoval et al. [27]. Besides the mentioned studies, others have focused on this type of vulnerability, but without exploring it in detail, as can be seen in [25, 29, 38, 40].
@@ V-A Uses of LLMs for Vulnerability Detection
A study by Zhang et al. [46] focused on LLMs capabilities for automated vulnerability detection. They tested over 10 different LLM models using various techniques like zero-shot and one-shot learning (for more details on prompting techniques, we refer the reader to Section VIII), and fine-tuning. They applied these models to various datasets of source code, exhibiting C/C++ and smart contract vulnerabilities. The study found that CodeLlama-7B achieved the highest F1-score of 82% with discriminative fine-tuning, along with a precision of approximately 89% and a recall of around 78%. In comparison, other methods were less effective.
According to Wen et al. [47], existing pre-trained models for vulnerability detection often miss important details and struggle with complex code. As a consequence, performing accurate vulnerability detection can be difficult. To address these challenges, Wen et al. [47] proposed a Structured Natural Language Comment tree-based Vulnerability Detection framework, based on pre-trained models, named SCALE. This framework is composed of three main modules. The first module, Comment Tree Construction, helps the model understand code semantics, by incorporating LLMs for comment generation and adding the comment node to Abstract Syntax Trees (ASTs). The second module, Structured Natural Language Comment Tree Construction, refines this process by explicitly involving code execution sequences and combining code syntax templates with the comment tree. The third module, SCT-Enhanced Representation, uses these improved comment trees to capture vulnerability patterns and better understand vulnerabilities in the code. Results show that SCALE outperforms existing methods, including supervised, retrained model-based, and other LLM-based approaches in detecting vulnerabilities.
@@ V-B Issues in Using LLMs for Vulnerability Detection
While LLMs show considerable promise for vulnerability detection, several challenges and limitations persist. In this section, we aim to unwrap the limitations highlighted by the literature. Table IV summarizes the challenges faced by LLMs in detecting vulnerabilities in code, as identified by the reviewed studies. In this table, we indicate with a green checkmark () which challenge was highlighted in which study. Notably, most studies focused on the prevalence of false positives, discussed in 7 papers, and challenges with complex tasks, highlighted in 6. In contrast, the other three challenges were mentioned in only one or two studies out of 12.
TABLE IV: Challenges in using LLMs for vulnerability detection identified in the studies
References False Positives Context Awareness Complex Tasks Verbosity and Non-Compliance Token Limits
Steenhoek et al. [52]
Happe et al. [53]
Yu et al. [54]
Ullah et al. [55]
Purba et al. [56]
Liu et al. [57]
Ding et al. [58]
Çetin et al. [59]
Pelofske et al. [60]
Chen et al. [61]
Sun et al. [62]
Gao et al. [63]
Zhang et al. [46]
Li et al. [64]
Hu et al. [65]
Zhou et al. [66]
Tamberg et al. [67]
Total number of studies 11 3 6 1 2
@@
False positives are one of the key challenges for LLMs, where the model incorrectly flags a non-existent vulnerability.
Various works highlighted the challenge faced by LLMs with false positives, when dealing with vulnerability detection. Ding et al. [58] reported that commonly used benchmarks often overestimate model performance, with a state-of-the-art 7B model achieving a 68.26% F1 score on a popular dataset BigVul, but only a 3.09% on the dataset they proposed. Similarly, Çetin et al. [59] found that GPT-4, despite being the top-performing model in their study, still had a false positive rate of 63%, with other models performing even worse, reaching up to 97%. Pelofske et al. [60] revealed that, while some models achieved high recall and precision rates (up to 100%) for certain vulnerability types, they still exhibited significant rates of false positives. Chen et al. [61] also noted that while ChatGPT was able to identify vulnerabilities in blockchain smart contracts, it struggled with precision, leading to false positives. Zhou et al. [66], Tamberg et al. [67], and Purba et al. [56] found that, although LLMs outperform traditional vulnerability detection approaches in identifying vulnerabilities, they also exhibit high false positive rates. We further discuss this aspect in Subsection V-C.
Some works have looked into proposing solutions to address this challenge. Li et al. [64] noticed that LLMs often produce false positives when used as static code analyzers. To address this, they proposed a method that prompts ChatGPT to summarize functions, breaking down tasks step by step and progressively refining results to improve accuracy. Their approach successfully eliminated 16 out of 20 false positives. Hu et al. [65] addressed the false positives issue by proposing a framework called GPTLens, which splits the conventional one-stage detection into two phases: generation and discrimination. In this framework, an LLM plays a dual role. In the generation phase, it acts as the Auditor, aiming to generate a wide range of vulnerabilities to capture the correct ones. In the discrimination phase, it acts as the Critic, evaluating these vulnerabilities to reduce false positives. Empirical results show that GPTLens significantly improves vulnerability detection over traditional methods.
Context awareness is another challenge that LLMs face when detecting vulnerabilities in code. Happe et al. [53] pointed out that hallucinations stem from a lack of contextual understanding and grounding. Sun et al. [62] addressed this issue by developing GPTScan, which enhances context awareness by combining GPT with static analysis techniques. GPTScan parses code, identifies potential vulnerabilities, and maps them to known types, significantly improving detection precision and recall.
Some researchers have reported that LLMs often struggle with more complex tasks, which often include real-world scenarios. For instance, Yu et al. [54] found that while GPT-4 outperformed other models in security code review, its performance was still limited in more complex scenarios. Ullah et al. [55] also highlighted that LLMs, including advanced models like GPT-4, exhibited unpredictable responses and incorrect reasoning when tasked with identifying security vulnerabilities in real-world scenarios. Moreover, Purba et al. [56] noted that LLMs performed well in detecting certain vulnerability patterns, but still could not surpass traditional tools in handling complex tasks, such as SQL injection and buffer overflow detection. Liu et al. [57] developed a benchmark to assess LLMs on progressively difficult tasks. They revealed that LLMs perform well (over 80% accuracy) at simpler tasks, such as identifying and classifying vulnerabilities. However, their accuracy dropped (below 30%) when faced with more complex tasks, such as detecting the root causes of vulnerabilities across extensive codebases. Finally, Ding et al. [58] observed that LLMs are not able to meet the requirements of effective vulnerability detection in real-world scenarios.
Yu et al. [54] found that, while GPT-4 was effective in detecting security vulnerabilities and outperformed other models, it faced verbosity and non-compliance challenges. These issues emerge when the model occasionally provides overly detailed responses or fails to meet prompt requirements, respectively. Both issues can affect the quality and trustworthiness of responses, in general.
Chen et al. [61] discussed the challenge of token limits, when ChatGPT is used for identifying vulnerabilities in blockchain smart contracts. They found that the maximum code sizes that GPT-4 and GPT-3.5 models could effectively analyze are around 33 KB and 16 KB, respectively. To address similar challenges posed by input length limitations, Zhang et al. [46] proposed a solution based on sliding windows that divide long sequences into overlapping sequences of tokens. In addition, they introduced a right-forward embedding strategy, which enhances context by making each token aware of tokens that come after.
Despite the challenges we have highlighted so far, most of the studies we have covered agree that LLMs have the potential to become effective tools for static analysis, provided additional fine-tuning and context-aware enhancements.
V-C Comparison of LLMs with Traditional Vulnerability Detection Approaches